OpenStudy (anonymous):

use the function rule f (x) = 3x - 2 find the output f (1.5) 2.5 -2.5 6.5 -6.5

OpenStudy (anonymous):

First plug in f(1.5) which is the same thing as f(3/2). f(3/2) = 3(3/2) -2 = 9/2 -2 multiply the -2 by 2 on the top and bottom to give it the same denominator f(3/2) = 9/2 - 4/2 = 5/2 which is f(1.5) = 2.5

OpenStudy (anonymous):

thank you! : )

OpenStudy (anonymous):

@ajmayberry hey can you help me with one last question?